### v2

# RE-RAG: Improving Open-Domain QA Performance and Interpretability with Relevance Estimator in Retrieval-Augmented Generation

Abstract
The Retrieval Augmented Generation (RAG) framework utilizes a combination of parametric knowledge and external knowledge to demonstrate state-of-the-art performance on open-domain question answering tasks. However, the RAG framework suffers from performance degradation when the query is accompanied by irrelevant contexts. In this work, we propose the RE-RAG framework, which introduces a relevance estimator (RE) that not only provides relative relevance between contexts as previous rerankers did, but also provides confidence, which can be used to classify whether given context is useful for answering the given question. We propose a weakly supervised method for training the RE simply utilizing question-answer data without any labels for correct contexts. We show that RE trained with a small generator (sLM) can not only improve the sLM fine-tuned together with RE but also improve previously unreferenced large language models (LLMs). Furthermore, we investigate new decoding strategies that utilize the proposed confidence measured by RE such as choosing to let the user know that it is “unanswerable” to answer the question given the retrieved contexts or choosing to rely on LLM’s parametric knowledge rather than unrelated contexts.
[…]
RE-RAG: Improving Open-Domain QA Performance and Interpretability with Relevance Estimator in Retrieval-Augmented Generation
Kiseung Kim Jay-Yoon Lee††thanks: Corresponding author Graduate School of Data Science, Seoul National University {kkskp, lee.jayyoon}@snu.ac.kr
1 Introduction
In recent years, the retrieval augmented generation framework has shown promising progress in natural language generation, specifically on knowledge-intensive tasks. This approach has been studied in many forms, from traditional RAG Lewis et al. (2020b), which aggregates answers from multiple contexts using document relevance scores as weights, to approaches like RALM Ram et al. (2023), which simply utilizes concatenated context as an in-context learning approach for large-language models (LLMs). Retrieval augmented generation enhances the model’s faithfulness and reliability by leveraging nonparametric knowledge on top of parametric knowledge Luo et al. (2023). In particular, the RAG framework has the advantage of being easily adaptable to modern LLMs Brown et al. (2020); Touvron et al. (2023). These advantages have sparked a significant amount of new research Asai et al. (2023); Lin et al. (2023); Shi et al. (2023) focused on the RAG framework.
Despite the great potential of the retrieval augmented generation framework, if the language model is provided with contexts that are not relevant to the query, it will be distracted by these inappropriate contexts, negatively affecting the accuracy of the answers Yoran et al. (2023). While retrievers or re-rankers in existing research have been effective at measuring the relative ranking across contexts to a query, these modules often fail to determine whether top-ranked contexts are actually relevant to the query or not. Furthermore, if a precise relevance score is not used in the traditional RAG framework, it can cause problems such as directing attention to documents that are less likely to answer the query.
In this work, we propose the RE-RAG framework, which extends traditional RAG by incorporating a relevance estimator (RE) to simultaneously measure the precise relative relevance between retrieved contexts and evaluate their confidence, which can be used to classify whether given context is useful for answering the given question. By more accurately measuring the relative relevance between contexts, RE computes precise relevance scores suitable for weighted aggregated answers in the traditional RAG framework and also acts as an efficient reranker. RE trained on a small generator (sLM) not only benefits sLM fine-tuned together with RE but can also be separated and applied to LLMs as well, benefiting both.
By explicitly classifying whether the context is useful for answering the query, the confidence of context measured by RE provides various decoding strategies. If the retrieved context set is irrelevant, we can choose to classify the query as “unanswerable”, while maintaining most of the accuracy for the answerable set. Additionally, if a low-confidence context set is retrieved, which will likely result in wrong answers by parroting the context as is Jia and Liang (2017), we can instead selectively leverage the LLM’s parametric knowledge to improve answer accuracy in most cases.
The main contributions of our work are:
1.
We propose a new framework called RE-RAG by adding an external Relevance Estimator (RE) module. We further suggest a weak supervision training method that can train RE without explicit labeled data on question-context compatibility. (§2.2)
2.
We demonstrate that RE-RAG, enhanced with RE, significantly improves upon the existing RAG. Addtionally, we show that RE trained on a small language model can improve the answer performance of LLMs. (§4.1)
3.
We propose to use the confidence level of the context set measured by RE to answer “unanswerable” for unanswerable context sets with minimal negative effects, or to complement LLM’s parametric knowledge. (§5.1)
2 Method
In this section, after reviewing the traditional RAG framework, we present the RE-RAG model combined with our relevance estimator.
2.1 Traditional RAG overview
Retriever Retriever searches for information in an external knowledge base and returns a related context set . In general, RAG systems use a bi-encoder type retriever such as DPR (Karpukhin et al., 2020), which is effective and fast in retrieving information. A question and a context are input to the encoder independently to obtain an embedding of , . The similarity score is calculated from the obtained embedding and then used to perform top- context retrieval.
Generator Generators that utilize the sequence-to-sequence model typically take a question and context as input and produce an answer with probability .
Answer marginalization RAG (Lewis et al., 2020b) introduced the answer generation models of RAG-sequence and RAG-token. We focus on the RAG-sequence model which marginalizes probability of where is an aggregated set of . which achieves higher performance than the RAG-token model and ensures the interpretability of the answer generation process. Individually generated answers per are marginalized as using the similarity score as shown in eq.(2).
2.2 RE-RAG framework
The retriever similarity score is trained to achieve high recall when retrieving multiple contexts, however, it was not initially designed to provide fine-grained relevancy score for aiding RAG generation steps in eq.(2). To address this issue, we propose a relevance estimator (RE) that re-ranks contexts and provides precise relevance scores to the generator.
Relevance Estimator
Relevance estimator (RE) measures the relevance between a question and context. We utilize a similar architecture to Nogueira et al. (2020) which utilizes a sequence-to-sequence model as a passage reranker.
Our RE receives the same input of question and context as the generator, but is trained to generate a classification token ("true" or "false") based on the relevance of the context to the input question. We normalize the probability of generating "true" and "false" tokens to get the final probability of generating the classification token. The obtained probability of a "true" token can independently be an indicator of the relevance of a single context to a given question. When comparing between multiple contexts, the "true" token probability can be converted to logit and used as the relevance score of the retrieved context.
Reranking of contexts by relevance
With the trained relevance estimator RE, we can rerank contexts in the initial retrieved set by their relevance and only take top- contexts to redefine before the answer-generation step. With a precise relevance score from RE, we can expect the RE-RAG to be more efficient, i.e. stronger performance with lower computation (see §4.2).
Answer marginalization with context RE
The question and context are concatenated and input to the generator model, and the generator generates per question. We replace the probability distribution in eq.(2) with the relevance scores from context RE to form eq.(6) as following:
We can expect higher performance with the marginalized answer if RE can provide an accurate relevance distribution (see §5.2).
2.3 Joint training of RE-RAG
We propose to utilize three different types of losses to train RE-RAG with our proposed relevance estimator. First, to train the generator model, we use a loss that combines the commonly used negative likelihood loss for ground truth with a probability that represents the relevance of the question and context.
simultaneously adjusts the probability of generating the classification token for the relevance estimator while training the generator.
Second, to obtain a learning signal for training the relevance estimator, we calculate the log-likelihood loss of the generator per retrieved context and compute its distribution across contexts as follows:
The log-likelihood loss varies depending on whether an answer can be inferred from the input context. Therefore, applying the softmax function to the log-likelihood loss values yields a probability distribution that represents the relevance between the given set of contexts and the question. We do not leverage any labeled data that entails the relevance of questions and contexts.
represents relative relevance between and
We calculate the KL-divergence loss between the probability distributions of the generator and the RE, and use this loss to train the model.
Lastly, in addition to applying a training loss on the probability of generating the classification token, we need to set an additional loss to prevent the RE from generating tokens other than the classification token. To do this, we utilize the additional loss as the sum of the probability of RE of generating all tokens other than classification token.
To train an effective system, the two models are trained jointly utilizing all three losses as follows:
where and are hyperparameters that act as scaling factors to balance the impact of each loss.
| Model | Extra | Generator | NQ | TQA | # Contexts |
| Small language models () | |||||
| RAG (Lewis et al., 2020b) | - | 445M | 44.5 | 56.8 | 50 |
| (Izacard and Grave, 2021b) | - | 220M | 48.2 | 65.0 | 100 |
| (Izacard and Grave, 2021b) | - | 770M | 51.4 | 67.6 | 100 |
| (Izacard and Grave, 2021a) | - | 220M | 50.1 | 69.3 | 100 |
| (Izacard and Grave, 2021a) | - | 770M | 54.4 | 72.5 | 100 |
| ReAtt (Jiang et al., 2022) | - | 770M | 54.7 | - | 100 |
| (Izacard and Grave, 2021a) | - | 220M | 48.6 | 67.4 | 25 |
| (Izacard and Grave, 2021a) | - | 770M | 53.9 | 71.2 | 25 |
| R2-D2 (Fajcik et al., 2021) | 125M | 1.04B | 55.9 | 69.9 | 25 |
| 220M | 220M | 49.9 | 68.2 | 25 | |
| 770M | 220M | 51.4 | 69.5 | 25 | |
| 770M | 770M | 54.0 | 70.2 | 25 | |
| Large language models () | |||||
| (Asai et al., 2023) | - | 7B | - | 66.4 | 5 |
| (Asai et al., 2023) | - | 13B | - | 69.3 | 5 |
| 770M | 7B | 45.7 | 67.1 | 5 | |
| 770M | 13B | 46.6 | 70.8 | 5 | |
| RA-DIT (Lin et al., 2023) | - | 65B | 43.9 | 75.1 | 10 |
| - | 70B | 38.1(40.7) | 63.5(66.3) | 10 | |
| - | 70B | 39.5(46.8) | 68.1(72.1) | 10 | |
| - | 175B | 42.9(45.9) | 69.0(70.7) | 10 | |
| (Shi et al., 2023) | - | 175B | 45.5 | 77.3 | 10 |
| 770M | 8B | 49.6 | 73.0 | 10 | |
| 770M | 70B | 48.0 | 72.4 | 10 | |
| 770M | 70B | 50.8 | 75.5 | 10 | |
| 770M | 175B | 49.3 | 72.6 | 10 | |
3 Experimental Setup
We evaluated the performance of our model on an open-domain QA dataset. In this section, we describe the dataset we used in our experiments and the details of our experiments.
3.1 Dataset
We evaluate our performance on two open-domain QA datasets:Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017). To train and evaluate our model, we utilize the context datasets retrieved for each question from NQ and TQA, as used in FiD-KD (Izacard and Grave, 2021a) and Akari (Asai et al., 2022). The dataset includes the top-20 training contexts, while the dev and test sets contain the top-100 contexts retrieved by the retriever. We used 20 contexts for training and the top-25 contexts extracted by the RE from the top-100 retrieved contexts for inference.
Natural Questions Natural Questions (Kwiatkowski et al., 2019) is a dataset of real questions asked by users on the web. The dataset consists of questions collected from the web, a long answer that can be viewed as gold context for the question, and a short answer with a short span. The open-domain QA version dataset of Natural Questions is a dataset that collects only questions where the answer span of the short answer is 5 tokens or less in length. We use the NQ-open dataset.
TriviaQA TriviaQA (Joshi et al., 2017) is a dataset of question-answer pairs collected from trivia enthusiasts. Each question and answer in the dataset has been reviewed by human annotators. We want to use the unfiltered version of TriviaQA dataset.
3.2 Evaluation Metric
The predicted answers are evaluated using EM score, a commonly used metric as in Izacard and Grave (2021b), Rajpurkar et al. (2016). The generated answers are normalized (e.g., lowercase, punctuation, article stripping) and compared to the correct answers in the dataset. We consider a generated answer to be correct if it exactly matches one of the correct answers in the given dataset after normalization.
3.3 Baseline
We investigate whether the performance of RE-RAG is competitive with that of the FiD (Izacard and Grave, 2021b)-based system. FiD has achieved excellent performance on the Question-Answering task, and the FiD-based application system also outperforms the RAG (Lewis et al., 2020b)-based system on the QA task.
We consider an additional baseline to compare the performance of RE when applied to LLMs. We compare the performance of RE and FiD-KD retriever when applied to LLMs. When applying the FiD-KD retriever to LLMs, we compared two methods: traditional RAG, which uses the retriever similarity score to perform answer marginalization, and RALM, which concatenates all context. Furthermore, we compare our performance with other studies Asai et al. (2023); Lin et al. (2023); Shi et al. (2023) that have implemented RAG in LLMs.
3.4 Model
The two components of our framework, RE and the generator, utilize the T5 model (Raffel et al., 2020). We utilize the T5-base, T5-large models, and explore three different model sizes depending on the combination of the two models.
Additionally, we utilize Llama2 (7B, 13B, 70B), Llama3 111https://github.com/meta-llama/llama3 (8B, 70B), and ChatGPT (“gpt-3.5-turbo-0125” version) as generators to assess if RE brings performance improvements when applied to LLMs. In our experiments, the LLMs used as generators are not fine-tuned for the downstream task.
4 Experiment Results
We investigate the QA performance of the RAG system with our newly proposed relevance estimator (RE). In addition to the QA performance of the whole system, we also examine the performance of the RE independently.
4.1 Main Results
The overall accuracy of our system on the two datasets (NQ and TQA) is shown in Table 1. Compared to the traditional RAG, our system, RE-RAG, performs better despite having the same total number of parameters. Our proposed RE improves the reliability of the RAG system by more accurately measuring the relevance between question and context. Our model performed competitively with models based on FiD structures(Izacard and Grave, 2021a; Jiang et al., 2022; Fajcik et al., 2021).
The accuracy of the RE module when applied to Large Language Models (LLMs) is shown at the bottom of Table 1. We only included the RAG-based model in our comparison because the FiD-based model is not applicable to LLMs due to structural differences. The RE module outperforms the FiD-KD retriever when applied to LLMs. When the RE module is applied to Llama2, it surpasses the Self-RAG, where the LMs themselves inspect the retrieved context and generated answers. In TQA, REPLUG with Codex scores slightly higher. The performance of TQA seems to depend more on the generator model than NQ (see Figure 2 for a related discussion), and we believe that this is the reason for the performance difference with Codex. Our model performs better on NQ, which is a more knowledge intensive task.
| Dataset | Model | Recall@k | |||
| R@1 | R@5 | R@10 | R@20 | ||
| NQ | FiD-KD | 49.4 | 73.8 | 79.6 | 84.3 |
| 46.2 | 72.4 | 80.1 | 84.7 | ||
| 59.5 | 77.8 | 82.7 | 85.5 | ||
| 61.9 | 79.4 | 83.6 | 86.4 | ||
| TQA | FiD-KD | 60.1 | 77.0 | 80.9 | 83.6 |
| 64.7 | 79.7 | 82.9 | 84.8 | ||
| 67.0 | 81.5 | 83.6 | 85.4 | ||
| 70.4 | 82.2 | 84.4 | 86.1 | ||
| Dataset | Model | Recall | Precision | F1 |
| NQ | FiD-KD | 73.2 | 21.9 | 33.7 |
| 10.3 | 31.0 | 15.5 | ||
| 51.3 | 33.9 | 40.9 | ||
| 45.9 | 38.3 | 41.7 | ||
| TQA | FiD-KD | 64.3 | 24.5 | 35.5 |
| 27.2 | 34.2 | 30.3 | ||
| 38.9 | 46.7 | 42.5 | ||
| 39.0 | 43.2 | 41.0 |
4.2 Performance of RE as a reranker and unanswerable set classifier
Table 2 shows the performance of our proposed RE-RAG’s RE as a reranker. For the Recall@k metric, we use the retrieval accuracy used by DPR (Karpukhin et al., 2020), FiD-KD (Izacard and Grave, 2021a), and ColbertQA (Khattab et al., 2021). Although the comparison retriever has been enhanced through knowledge distillation methods using FiD attention scores, our proposed RE still demonstrated superior performance. In particular, RE performs better as the number of contexts decreases, which means that RE is more efficient when there are fewer contexts to utilize.
Table 3 shows the performance of the context relevance estimator (RE) as a “unanswerable” set classifier. “unanswerable” set means that the context set of the top-25 contexts does not contain a gold answer in any context. For classification, we used the cosine similarity score of the hidden representation of the question and context for retriever and the probability of generating a “true” token by the model for RE and MonoT5 (Nogueira et al., 2020). For the optimal threshold, we searched for the value that maximizes F1 score in steps of 0.1 from 0.5 to 0.9 at dev set.
Our RE showed better “unanswerable” set classification performance than FiD-KD retriever or MonoT5 based on F1 score. Looking at the detailed performance, we found that the retriever performed better for recall, but the RE performed better for precision. This is because the retriever classified a large number of context sets as all “unanswerable” sets, while our proposed RE showed a good balance between precision and recall.
| Dataset | Model | Score | Answerable context set | |
| O | X | |||
| NQ | FiD-KD | 58.3 32.7 | 73.4 | |
| RE | 58.3 54.9 | 51.3 | ||
| FiD-KD | 61.5 34.9 | 71.3 | ||
| RE | 61.5 57.9 | 45.9 | ||
| TQA | FiD-KD | 78.7 51.2 | 63.5 | |
| RE | 78.7 77.0 | 38.9 | ||
| FiD-KD | 80.4 51.6 | 62.7 | ||
| RE | 80.4 77.9 | 39.0 | ||
5 Analysis
5.1 Exploring decoding strategies in low confidence context sets
In this section, we review two strategies that can be used when a context set with a low confidence score is retrieved. The confidence score for a context set is determined using the maximum value of the “true” token probability computed by RE for the contexts within the set. We examine the strategy of answering “unanswerable” when a low confidence context set is returned in a small Language Model (sLM), where parametric knowledge is scarce. Additionally, we examine the strategy of directly utilizing parametric knowledge in Large Language Models (LLMs), where parametric knowledge is abundant.
Classify as “unanswerable” Table 4 shows the change in accuracy after letting the model respond with “unanswerable” when the retrieved context set has low confidence. For the confidence threshold value that determines whether the model should respond with “unanswerable”, we chose the value that optimizes the classification performance as determined in Table 3. We evaluate the accuracy by dividing the entire test set into answerable sets, which contain at least one gold answer in the context set, and unanswerable sets, which contain none.
Our RE model shows relatively minor accuracy loss on the answerable set when responding with “unanswerable” for context sets measured with low confidence, but gains significant ability on the unanswerable set. In contrast, the FiD-KD retriever loses a substantial amount of accuracy on the answerable set when it responds with “unanswerable” for low-confidence context sets, resulting in a larger negative effect compared to our model. If we want to preserve the answerable set accuracy of the FiD-KD retriever, its ability to classify “unanswerable” is significantly reduced compared to RE (see Appendix E).
| P-Generator | R-Generator | NQ | TQA |
| (NQ: 31.1/TQA: 64.3) | 46.2 45.9(-0.3) | 68.0 69.3(+1.3) | |
| 47.3 46.5(-0.8) | 71.5 72.1(+0.6) | ||
| 48.0 46.9(-1.1) | 72.4 72.9(+0.5) | ||
| (NQ: 41.3/TQA: 75.1) | 49.6 49.8(+0.2) | 73.0 75.4(+2.4) | |
| 50.8 50.8(-) | 75.5 76.7(+1.2) | ||
| ChatGPT (NQ: 37.7/TQA: 72.0) | ChatGPT | 49.3 49.3(-) | 72.6 73.6(+1.0) |
Selectively using parametric knowledge We further explore how we can effectively utilize the rich parametric knowledge of LLMs. When the confidence of the retrieved context is low, we examine a mixed strategy that optionally bypasses the context and relies solely on the parametric knowledge of the largest model to generate the correct answer. For the confidence threshold value that determines whether the model should answer using only parametric knowledge, we selected the value that optimizes classification performance as determined in Table 3. For each type of model, we utilize the one with the largest number of parameters as the parametric knowledge base.
Table 5 shows the change in accuracy when decoding the answer using the mixed strategy. In most cases, our strategy achieves accuracy gains in TQA without significant losses in NQ, except in cases where parametric knowledge is particularly scarce, such as in NQ on Llama2. NQ is a more knowledge-intensive task compared to TQA, where there is less benefit from utilizing parametric knowledge.
When parametric knowledge can be used effectively, the mixed strategy achieves larger gains in smaller models, and the performance gap narrows compared to larger models. Figure 2 illustrates the relationship between confidence score and accuracy by model size. At high confidence scores on the TQA dataset, small size models achieve similar accuracy to large size models. At low confidence scores, the difference in performance between small and large models becomes more pronounced. When using small size models, higher efficiency can be achieved by utilizing retrieval augmented generation only when a high confidence context set is retrieved, and selectively leveraging the parametric knowledge of large size models when a low confidence context set is retrieved.
| Model | NQ | TQA |
| Baseline | 39.5 | 54.9 |
| Baseline w/ RE score | 43.1 | 60.1 |
| Baseline w/ RE rerank | 46.8 | 63.9 |
| Baseline w/ RE rerank, score | 49.6 | 67.8 |
| 49.9 | 68.2 |
5.2 Ablation Study
Effectiveness of RE We perform an ablation study to investigate the effectiveness of the added RE in RE-RAG. The effect of our proposed RE is twofold. First, it performs better re-ranking than the retriever, selecting more accurate context and passing it to the generator. Second, it calculates a more accurate relevance score than retriever’s similarity score and uses it in the answer marginalization process. In Table 6, the performance of methods with each component of the RE added is presented, using a model that was trained with only the T5-base generator, after removing the RE, as the baseline.
We construct the following experiment to isolate the two effects. First, we apply the top 25 contexts from retriever and their similarity scores to the baseline model. Next, there are the top-25 contexts from the retriever with the RE’s score applied (RE score) and the top-25 contexts from the RE with the retriever’s similarity score applied (RE rerank). Finally, we compare the performance of applying the RE’s top-25 contexts and score to the baseline model (RE rerank, score).
Both effects of the RE are found to be significant in improving the performance of the baseline model. This shows that not only the quality of the context input to the generator plays an important role, but also the score, which means the importance of each context.
| Model | NQ | TQA |
| Baseline | 0.435 | 0.561 |
| - normalization | 0.0005 | 0.0002 |
Remove training components We investigate the impact of removing the regularization process in eq.(3) on the classification performance of RE while training on the model. Table 7 shows how the “true” token probability level output by the RE changes when the normalization process is removed. It can be seen that when the normalization process is removed, RE can only perform the function of re-ranking but loses the function of measuring confidence. This is because the normalization process allows the model to adjust its output strictly between “true” and “false” tokens.
Table 8 shows the difference in EM scores on the dev set when is removed from the training process. We observed that removing from the training process decreases answer performance. We believe that contributes to achieving more optimal performance by using loss information from generator to directly propagate the relative importance of contexts to the RE.
6 Related Works
Previous research has shown that the performance of Question Answering systems can be improved by utilizing external knowledge about questions (Chen et al., 2017). Methods for more accurate retrieval of external knowledge Karpukhin et al. (2020); Khattab et al. (2021); Gao and Callan (2022) have been studied to make these systems more efficient. In open-domain QA, models that extract and use answers from retrieved documents have been studied Karpukhin et al. (2020); Khattab et al. (2021); Cheng et al. (2021), but studies that utilize generative models such as T5 (Raffel et al., 2020) or BART (Lewis et al., 2020a) have become more common Lewis et al. (2020b); Izacard and Grave (2021b). RAG and FiD achieved powerful performance in open-domain QA using different methods. Subsequently, models Izacard and Grave (2021a); Fajcik et al. (2021); Singh et al. (2021); Jiang et al. (2022) that leverage and improve upon the structural advantages of FiD have been proposed. For Atlas (Izacard et al., 2022), state-of-the-art performance was achieved through an improved retriever (Izacard et al., 2021) and scaling up the model. In the case of RAG, there is a study that improved performance by introducing a BERT (Devlin et al., 2019)-based reranker (Glass et al., 2022), but it utilized additional data and high-quality label data when training the reranker.
Recently, large language models (LLMs) such as GPT (Brown et al., 2020) and Llama (Touvron et al., 2023), which have been developed in recent years, face limitations with FiD methods that require encoded data. Consequently, research on RAG models, which can directly input context, has received renewed attention. Asai et al. (2023); Lin et al. (2023); Shi et al. (2023) These approaches have achieved performance improvements by training a retriever, which can also be applied to LLMs, or by performing the review of questions and context within the model itself.
| Model | NQ | TQA |
| Baseline | 49.1 | 67.8 |
| - | 48.0 | 66.7 |
7 Conclusion
We propose the RE-RAG framework, which extends traditional RAG by incorporating RE that can measure the relative relevance and confidence of contexts. We demonstrate that the RE-RAG framework can enhance the performance of traditional RAG. We show that the RE module, as a detachable component, can be combined with modern large language models (LLMs) to improve their performance. Furthermore, we exploree some decoding strategies that leverage the confidence information measured by the RE module to either answer “unanswerable” or selectively utilize the parametric knowledge of the LLMs when a low confidence context set is retrieved. We hope that our research will inspire the exploration of various additional modules for retrieval-augmented generation.
8 Limitation
Our research has primarily focused on improving answer performance in single-hop QA tasks. We have not sufficiently verified the effectiveness of our proposed framework in multi-hop QA tasks. We believe that in the future, we can explore whether the RE-RAG framework can be extended to multi-hop QA.
In our work, we explored a decoding strategy that measures with confidence whether a context is truly useful for a query and classifies low confidence contexts as unanswerable. However, a truly unanswerable query is one where the query cannot be adequately answered even when utilizing the model’s parametric knowledge. We believe that future research needs to be conducted to detect whether the parametric knowledge has knowledge that can adequately answer the query in order to finally classify the unanswerable problem.
Appendix A Dataset Statistics
Table 8 shows the statistics for the Natural Questions and TriviaQA unfilitated datasets we used.
| Dataset | Train | Dev | Test |
| Natural Questions | 79,168 | 8,757 | 3,610 |
| TriviaQA | 78,785 | 8,837 | 11,313 |
| Question | Context | Gold Answer | "True" prob |
| who played mark on the show the rifleman | …Mark McCain is the son of fictitious rancher Lucas McCain in the ABC Western television series "The Rifleman," starring Chuck Connors, which ran from 1958 to 1963. Singer/actor and former Mouseketeer Johnny Crawford was cast in the role and… | John Ernest Crawford | 0.987 |
| when does the cannes film festival take place | …2017 Cannes Film Festival The 70th Cannes Film Festival took place from 17 to 28 May 2017, in Cannes, France … | Cannes, France, usually in May | 0.994 |
| how many strong verbs are there in german | …Germanic strong verbs are commonly divided into 7 classes, based on the type of vowel alternation. This is in turn based mostly… | more than 200, more than 200 strong | 0.949 |
| how many episodes of corrie has there been | …The show airs six times a week: Monday, Wednesday and Friday 7:30-8 pm and 8:30-9 pm. Since 2017, ten sequential classic episodes of the series from 1986… | 9,436 | 0.147 |
Appendix B Training Details
We used T5-base with a parameter size of 223M and T5-large model with a parameter size of 770M as modulators in all experiments. We trained the system on 4 A6000 GPUs, while and were trained on 2 A100 and 4 A100 GPUs, respectively.
We used a constant learning rate of for all sizes of RE-RAG systems. We used AdamW as the optimizer and weight decay was . For batch size, we used gradient accumulation for all sizes of models, resulting in an effective batch size of 64. For the hyperparameters that balance the proposed losses, we utilized the default value of 1 for both and . We did not explore hyperparameters that achieve better performance due to time and limited computing resources.
For model selection, we evaluated every 1 epoch and selected the case with the highest answer accuracy of the dev set. The dev set answer accuracy was measured using the top-10 context of the RE. Since the answer accuracy of the top-10 context of the RE is similar to the answer accuracy of the top-25 context, this helped to save computational resources and time while still producing valid results.
| P-Generator | R-Generator | NQ | TQA |
| (N31.1/T64.3) | 36.1 35.8(-0.3) | 58.4 62.8(+4.4) | |
| 38.8 36.9(-1.9) | 64.9 65.4(+0.5) | ||
| 40.7 37.4(-3.3) | 66.3 66.2(-0.1) | ||
| (N41.3/T75.1) | 38.2 42.1(+3.9) | 57.6 66.9(+9.3) | |
| 46.8 45.6(-1.2) | 72.1 74.0(+1.9) | ||
| ChatGPT (N37.7/T72.0) | ChatGPT | 45.9 43.2(-2.7) | 70.7 72.1(+1.4) |
Appendix C Effectiveness of the RE
We perform a qualitative analysis to see if our proposed relevance estimator (RE) is effectively classifying relevant contexts. Table 3 shows a few contexts in the NQ test set.
Some of the contexts that the RE predicts are highly relevant to the question even when they do not contain the exact ground truth answer. The first few examples in Table 3 are examples that are categorized as true context because they contain phrases that are semantically equivalent to the correct answer albeit not having the exact same form in the context. This shows that although the RE is trained to measure the relevance of a question to a context through a limited set of ground truth answers, it is actually capable of measuring a broader range of relevance.
In addition to the examples above, there are cases where the RE misclassified contexts as containing the correct answer. As shown in the example in Table 10, the RE classified the context containing “the number of classes of strong verbs in German” as the correct context for the question about “the number of strong verbs in German”, which means that our RE is still limited in its ability to capture the fine-grained meaning of the question in the retrieved context. On the other hand, in the last example, for the question about “the number of episodes”, it succeeded in classifying the context containing “the number of classical episodes” as an incorrect context.
| Dataset | Type | Threshold | ||||
| 0.5 | 0.6 | 0.7 | 0.8 | 0.9 | ||
| NQ | Answerable | 61.3 | 56.2 | 34.9 | 6.4 | 0.0 |
| Unanswerable | 2.3 | 27.8 | 71.3 | 97.2 | 99.8 | |
| TQA | Answerable | 77.3 | 51.6 | 9.2 | 0.1 | 0.0 |
| Unanswerable | 14.3 | 62.7 | 94.7 | 100.0 | 100.0 | |
Appendix D Selectively using parametric knowledge with FiD-KD
Table 11 shows the change in EM score when applying the mixed decoding strategy, using the cosine similarity score of the FiD-KD retriever as the confidence score. For small parameter generators, the EM score is low when applying the FiD-KD retriever to LLMs, which results in a high gain when utilizing parametric knowledge of large parameter models. However, since the classification performance of the FiD-KD retriever is lower than that of RE, even utilizing parametric knowledge does not significantly outperform the baseline performance of parametric knowledge. Especially for more knowledge-intensive tasks such as NQ, the performance loss is substantial.
Appendix E FiD-KD retriever’s performance in “unanswerable” scenarios
Table 12 shows the performance of the FiD-KD retriever in unanswerable scenarios according to different threshold values. For the FiD-KD retriever, it is observed that while trying to maintain performance on the answerable set, the classification ability on the unanswerable set significantly decreases.